# Share the Sugar

### v1

@@ 2 Negative Dominance
Stochastic Dominance is extremely plausible, and among the most widely accepted principles in normative decision theory. (For an array of arguments in its favor, see Tarsney (2020).) It is compatible with a very wide range of attitudes toward risk, being satisfied not just by standard expected utility theory but also by more permissive alternatives like rank-dependent/risk-weighted expected utility (Quiggin, 1982; Buchak, 2013) and the version of weighted-linear utility theory recently defended by Bottomley and Williamson (forthcoming). [^9] To put the point more exactly, a classic result from Hadar and Russell (1969) and Hanoch and Levy (1969) implies that an expected utility maximizer will satisfy Stochastic Dominance with respect to a given ranking of outcomes if and only if her utility function is increasing with respect to that ranking, i.e., if and only if she prefers certainty of a better outcome to certainty of a worse outcome. The same thing turns out to be true of the alternatives to expected utility theory mentioned in the main text. So the fact that Statewise Negative Dominance and Stochastic Dominance conflict is a strong argument against Statewise Negative Dominance. But no similar argument can be given against (plain) Negative Dominance, since it is consistent with Stochastic Dominance (Lederman, 2023b, Proposition 3.1).
@@ 3 Personal Good
How should the relation of personal betterness for individuals connect to the relation of impersonal betterness? The following articulates a weak constraint on this connection:
@@
If the same individuals exist in outcomes $o$ and $o^{\prime}$, and $o$ is personally at least as good as $o^{\prime}$ for every individual and strictly personally better for at least one individual, then $o$ is impersonally better than $o^{\prime}$.
This principle only concerns outcomes; it doesn’t tell us anything about lotteries in general.
How might it be extended to lotteries as well as outcomes? We will work with the following weak extension of Pareto:
@@
Suppose two lotteries $L$ and $L^{\prime}$ are such that (1) every possible individual $i$ has the same probability of existence in $L$ as in $L^{\prime}$. Then, (i) if for every individual $i$ and life $l$, lottery $L_{1}$ gives $i$ at least as great a probability of a life at least as personally good as $l$, compared to $L_{2}$, then $L_{1}$ is at least as impersonally good as $L_{2}$; and (ii) if in addition, for some $i$ and $l$, $L_{1}$ gives $i$ a strictly greater probability of a life at least as personally good as $l$, then $L_{1}$ is strictly impersonally better than $L_{2}$. [^10] Our arguments below will rely mainly on part (ii) of Personal Good. We will appeal to part (i) only in §6.
@@
This principle extends Pareto to lotteries, and to cases where different numbers of people may exist in each outcome. [^11] McCarthy et al. (2020) advocate a closely related principle which they call “Anteriority”. The extension is very modest in both respects. For the first, it only provides comparisons where each individual has a fixed probability of existence. This means that it is compatible with a wide range of views about how to value nonexistence—when a choice affects any individual’s probability of existence, the principle remains silent. Personal Good also offers a modest extension in the way that it handles individual betterness for lotteries. In particular, it implicitly invokes a very weak sufficient condition for $L_{1}$ to be (weakly/strictly) better than $L_{2}$ from the perspective of an individual, namely, that the personal lottery (probability distribution over lives) she receives from $L_{1}$ (weakly/strictly) stochastically dominates the personal lottery she receives from $L_{2}$. As described above, Stochastic Dominance is a minimal and compelling principle for ranking lotteries, which is compatible with a very wide range of attitudes toward risk.
Personal Good should be especially compelling to those who take an “individualistic” view of population ethics, according to which our fundamental objects of concern should be the interests of each individual (rather than the impersonal value of the world as a whole), and so options should be compared in the first instance from the perspective of each affected individual. [^12] In associating Pareto-like principles with “individualism”, we are following Harsanyi (1955). From this perspective, it seems obvious that if $L_{1}$ is unambiguously better than $L_{2}$ from the perspective of some individuals, and equivalent from the perspective of everyone else, then it is better overall. Since Personal Good makes no commitments about the value of non-existence, it should also appeal to those who endorse “person-affecting” views in population ethics, which characteristically combine individualism with an unwillingness to compare existence and non-existence.
@@
With that said, and though we ourselves find the principle attractive, there is some precedent for rejecting it. Some egalitarians, for instance, will reject Personal Good because, while it allows concern for ex ante equality (equality between personal lotteries), it rules out any independent concern for ex post equality (equality between individual lives in a given outcome). We return to this idea, and in general, the possibility of rejecting Personal Good in §7.
@@ 4 Incomparable lives
We first make an assumption that we think of as basically a convenience, without significant conceptual import:
@@
For any finite set of possible individuals and any function from those individuals to lives other than non-existence, there is an outcome where exactly those individuals exist, with exactly those lives.
Moreover, for any finite set of outcomes and any probability distribution defined on that set, there is a lottery which assigns those probabilities to those outcomes.
We are assuming (as we said earlier) that the set of possible individuals is countably infinite, and that the impersonal betterness relation is defined over the whole set of lotteries. We won’t need anything like the full strength of Richness: all we need is that the domain of the betterness relation includes outcomes with the structure of our examples below.
@@
Suppose that outcomes $o$ and $o^{\prime}$ are identical except that one possible individual receives life $a$ in $o$ and life $b$ in $o^{\prime}$, where $a$ and $b$ are incomparable in terms of personal value. Then $o$ and $o^{\prime}$ are incomparable in terms of impersonal value.
As with Personal Good, this principle is a natural expression of “individualism” about impersonal betterness: We compare outcomes by comparing them from the vantage point of each individual, and when only one individual’s fate differs between two outcomes, the impersonal betterness ordering should reflect that individual’s personal betterness ordering. When the two lives that individual might receive are personally incomparable, their incomparability is “transmitted” to the outcomes in which they figure.
Finally, we will assume the following principle about the structure of incomparability on outcomes (which, to recall, we are treating as equivalent to lotteries which assign probability 1 to a single outcome).
@@
Suppose that outcome $o$ is impersonally incomparable with outcomes $o^{\prime}$ and $o^{\prime\prime}$. Then there is either
@@
    an outcome $o^{+}$ which is strictly better than $o$ for some individual, and equally good for everyone else, while still being impersonally incomparable with both $o^{\prime}$ and $o^{\prime\prime}$, or
@@
    an outcome $o^{-}$ which is strictly worse than $o$ for some individual, and equally good for everyone else, while still being impersonally incomparable with both $o^{\prime}$ and $o^{\prime\prime}$.
Insensitivity to small improvements or small worsenings is a defining feature of incomparability, by contrast to equal goodness. Thus it should not be controversial that, given incomparable $o$ and $o^{\prime}$, there is either an $o^{+}$ that improves $o$ and is still incomparable with $o^{\prime}$ or an $o^{-}$ that worsens $o$ and is still incomparable with $o^{\prime}$. The above principle adds to this uncontroversial claim two further ideas. First, it adds that we can achieve this improvement (worsening) by improving (worsening) the life of some individual in $o$. This is slightly stronger, but we do not expect it to be particularly controversial. Second, it adds that when $o$ is incomparable to two further outcomes $o^{\prime}$ and $o^{\prime\prime}$, there is a single improvement (or worsening) which preserves incomparability with both $o^{\prime}$ and $o^{\prime\prime}$. For instance, if an outcome where Patricia is an artist is incomparable with outcomes where she is a banker or a CIA operative, then we can either slightly improve or slightly worsen the first outcome (e.g., by slightly adjusting Patricia’s income) while preserving incomparability. Like the first, this claim seems plausible to us, and holds in simple, natural models of value incomparability. In any case, as described below, our argument makes only limited use of Small Improvements, and does not require it to hold in full generality.
@@
Proposition 4.1 (Individual-Level Incomparability).
Given Richness, the principles of Negative Dominance, Personal Good, Incomparability Transmission, and Small Improvements rule out incomparability in personal value.
This can be proved using a slight modification of the example from the introduction, illustrated in Table 2:
@@
Here again we are considering lotteries that affect just two possible individuals, $P_{1}$ and $P_{2}$. (We can assume that everyone else is assured of non-existence.) Given two personally incomparable lives $a$ and $b$, Incomparability Transmission implies that the outcomes $(a,b)$ and $(b,a)$ are both incomparable with the outcomes $(a,a)$ and $(b,b)$ (where the first and second terms in each ordered pair represent the lives of $P_{1}$ and $P_{2}$ respectively). Small Improvements implies that we can find an outcome which is exactly as good for one individual and strictly better for the other than $(a,a)$ while still being incomparable with both $(a,b)$ and $(b,a)$, or an outcome which is exactly as good for one individual and strictly worse for the other, while still being incomparable with both $(a,b)$ and $(b,a)$. Let’s concentrate on the first case, and assume the improvement is to the first individual, giving this person the life $a^{+}$. (The other cases are very similar.) The outcome $(a^{+},a)$ is by assumption still incomparable with $(a,b)$ and $(b,a)$. Richness then guarantees that the domain of the impersonal betterness relation includes the two lotteries described in Table 2. Both possible outcomes of Lottery 2 are incomparable with both possible outcomes of Lottery 1, so Negative Dominance implies that Lottery 2 is not better than Lottery 1. But Lottery 2 gives $P_{1}$ a stochastically dominant personal lottery, and $P_{2}$ a stochastically equivalent lottery, so Personal Good implies that Lottery 2 is strictly better.131313 Lederman (2023a, b) presents negative results showing a conflict between Negative Dominance and “unidimensional” principles that are analogous to but much stronger than our Personal Good. Throughout, he uses an assumption that in our setting would amount to taking personal betterness to be complete. The present result shows that, if personal betterness fails to be complete, then much weaker assumptions already suffice to generate an inconsistency. Similarly Lederman (2023b, Proposition 3.1) provides a consistency result for a package which includes Richness, Negative Dominance, and Personal Good. Our result shows that if personal betterness (or, in his setting, the “unidimensional” order) can be incomplete, then this result breaks under plausible further assumptions.
@@
This argument establishes, we think, that Negative Dominance and Personal Good are jointly incompatible with the existence of incomparable lives. To briefly consider the other ways out: Richness is only used to secure the existence of two lotteries with the structure of Lottery 1 and Lottery 2 in Table 2, which are hardly exotic. Incomparability Transmission is only used to secure the conclusion that, if $a$ and $b$ are personally incomparable, then $(a,a)$ and $(b,b)$ are impersonally incomparable with $(a,b)$ and $(b,a)$. And this claim seems very plausible in its own right, since if $a$ and $b$ are incomparable, what basis could we have for thinking that $(a,b)$ or $(b,a)$ are better or worse than $(a,a)$ or $(b,b)$? [^14] Given the possibility of small improvements, it is not plausible to claim that a pair of outcomes like $(a,a)$ and $(a,b)$ will in general be equally good. In our particular example, one might claim that $(a,a)$ is better than $(a,b)$ on grounds of equality, or alternatively that $(a,b)$ is better than $(a,a)$ on grounds of diversity. But these claims would not seriously undermine our argument. If you think that more diverse outcomes are better, for instance, this only seems to strengthen our claim that neither outcome of Lottery 2 is better than either outcome of Lottery 1. If you think that more equal outcomes are better, we could simply change the example to sweeten an outcome in Lottery 1 rather than Lottery 2. And finally, the application of Small Improvements should be uncontroversial: given that $(a,a)$ is incomparable with $(a,b)$ and $(b,a)$, then it seems (in the spirit of insensitivity to small improvements) that it should be possible to slightly improve or worsen someone’s life in $(a,a)$ while maintaining that incomparability. [^15] One might think that $(a^{+},a)$ is better than $(a,b)$ on the grounds that it is better for someone and worse for no one. But this reasoning, though intuitive, cannot be sound, since it would lead to cycles of betterness: for instance, given incomparable lives $a$ and $b$, we would get that $(a,a)$ is better than $(a^{-},b)$, which is better than $(b,b^{-})$, which is better than $(b^{-},a^{+})$, which is better than $(a,a)$. There is a close analogy here with the prima facie plausible decision-theoretic principle that if one gamble yields a better outcome than another in some state of nature, and does not yield a worse outcome in any state of nature, then it is better overall. While this looks like an unobjectionable statewise dominance principle, it leads to cycles in the presence of incomparability (MacAskill, 2013). The lesson is that Pareto (resp. statewise dominance) principles must require that the superior outcome (resp. gamble) is at least as good for every individual (resp. in every state), not just that it isn’t worse for any individual (resp. in any state). We think this principle is plausible in full generality, but, as with Richness, rejecting it in particular cases would not be enough to save the combination of Personal Good and Negative Dominance, since all our argument relies on is that there is some case with the structure of the one above. [^16] Nebel (2020) also presents an opaque sweetening-inspired puzzle in a population-ethical context. His puzzle compares the outcomes $(a,b)$ and $(b^{+},a^{+})$, where $a$ and $a^{+}$ are both incomparable with $b$ and $b^{+}$. He points out that the latter outcome is intuitively better than the former, and that this verdict can be derived from Pareto, anonymity, and transitivity, but that this conflicts with the “Person-Affecting Restriction” (PAR), which claims that one outcome is better than another only if there is someone for whom it’s better. Nebel sees this as a reason to give up PAR, and we agree. But giving up PAR does not resolve our puzzle, which uses only the much weaker and more plausible principle of Incomparability Transmission. (Unlike PAR, Incomparability Transmission is consistent with the other principles Nebel invokes, even in the presence of personal incomparability.) An alternative response to Nebel’s argument might be to reject his anonymity assumption, for instance weakening it to “relative anonymity” (which asserts that applying the same permutation of individual welfare levels to a pair of outcomes does not affect the evaluative comparison between them). But since we don’t rely on anonymity, this escape route likewise doesn’t generate a response to our argument. As Nebel does to Hare’s argument, it’s also possible to generate a structural analogue of our argument that inverts the roles of states and individuals. In this inverted argument, the analogue of Negative Dominance asserts that if no individual’s personal lottery under $L_{1}$ is at least as good as any individual’s personal lottery under $L_{2}$, then $L_{1}$ is not strictly better than $L_{2}$, while the analogue of Personal Good asserts that if in at least one state, the outcome of $L_{1}$ is Suppes-Sen superior to the outcome of $L_{2}$ (i.e., Pareto superior to a permutation of that outcome), and in all other states, the outcomes of $L_{1}$ and $L_{2}$ are identical, then $L_{1}$ is better than $L_{2}$. (To put these principles into conflict, just switch the rows and columns of Lottery 2 in Table 2, with $a^{+}$ in the top right. The argument will of course require auxiliary assumptions analogous to those used in Proposition 4.1.)
An important special case of the preceding result is where one of the incomparable “lives” is nonexistence. Many philosophers accept the following principle:
@@
There is some life $a$ such that, whenever two outcomes $o$ and $o^{\prime}$ are identical except that one possible individual does not exist in $o$ and exists with life $a$ in $o^{\prime}$, $o$ and $o^{\prime}$ are incomparable in terms of impersonal value.
One motivation for this principle is the belief that existence and non-existence are incomparable in terms of personal value, coupled with a principle like Incomparability Transmission. This might motivate the view that populations of different size are always incomparable (Bader, 2022). Alternatively, whatever one thinks about the personal value of nonexistence, one might accept Variable-Population Incompleteness to avoid undesirable conclusions about impersonal value. For instance, critical-range utilitarianism, which holds that adding an individual with slightly positive welfare to a population results in incomparability, can avoid both the Repugnant Conclusion (Parfit, 1984, §131, pp. 387–90) and the Sadistic Conclusion (Arrhenius, 2000). [^17] Gustafsson (2020) proposes a version of critical-range utilitarianism that is motivated by a fourth category of value called “undistinguishedness” and shows that it escapes versions of the Repugnant and Sadistic Conclusions. But the preceding argument shows us that both these views are in tension with Negative Dominance and Personal Good, taken together:
Proposition 4.2 (Variable-Population Incomparability).
@@
To see this, let $b$ in Table 2 represent nonexistence and $a$ be a life of the sort described by Variable-Population Incompleteness, such that adding someone to the population with life $a$ yields incomparability. Then Variable-Population Incompleteness implies that $(a,b)$ and $(b,a)$ are incomparable with $(a,a)$ and $(b,b)$. [^18] Note that, while the argument as stated appeals to the “empty population” $(b,b)$, we can avoid this by simply adding another person living life $c$ to all outcomes under consideration. The rest of the proof then follows the proof of Proposition 4.1. The upshot is that Negative Dominance and Personal Good together tell strongly against views that see changes in population size as a source of incomparability.

### v2

@@ 2 Negative Dominance
Stochastic Dominance is extremely plausible, and among the most widely accepted principles in normative decision theory. (For arguments in its favor, see Easwaran (2014: §3.4.4), (Russell forthcoming: §4), and Tarsney (forthcoming: §4).) It is compatible with a very wide range of attitudes toward risk, being satisfied not just by standard expected utility theory but also by more permissive alternatives like rank-dependent/risk-weighted expected utility (Quiggin 1982; Buchak 2013) and the version of weighted-linear utility theory recently defended by Bottomley and Williamson (2024). [^12] To put the point more exactly, a classic result from Hadar and Russell (1969) and Hanoch and Levy (1969) implies that an expected utility maximizer will satisfy Stochastic Dominance with respect to a given ranking of outcomes if her utility function represents that ranking, i.e., assigns greater utility to one outcome than another iff the first outcome is higher in the ranking. The same thing turns out to be true of the alternatives to expected utility theory mentioned in the main text. The fact that Statewise Negative Dominance is incompatible with Stochastic Dominance in the presence of incomparability therefore substantially undercuts the prima facie plausibility of Statewise Negative Dominance for anyone not already committed to completeness.
Some will nevertheless respond to this argument by giving up Stochastic Dominance in favor of Statewise Negative Dominance. Others, like us, will take it to show that anyone who believes in evaluative incomparability has good reason to reject Statewise Negative Dominance. But in either case, crucially, the preceding argument does not give any reason to reject Negative Dominance. First, Negative Dominance is consistent with Stochastic Dominance even in the presence of incompleteness (Lederman 2023: Proposition 3.1; Lederman forthcoming: §7). Second, in our view, Negative Dominance is more compellingly motivated than Statewise Negative Dominance. In particular, our earlier positive argument for Negative Dominance, which turned on grounding comparisons between lotteries in the final value of their potential outcomes, does not support Statewise Negative Dominance and is therefore not undermined by its rejection. Statewise Negative Dominance insists that comparisons between lotteries must be grounded in comparisons between their outcomes in the same state of nature. But nothing in our argument for Negative Dominance motivates this restriction. [^13] This claim may seem surprising given that Schoenfield (2014) and Doody (2023) have suggested that Statewise Negative Dominance can be motivated by a primary concern for the value of outcomes. But this concern does not justify an insistence on statewise comparisons. Instead, this insistence represents (in our view) a mistaken concern with counterfactuals—with comparing the actual outcome of one’s choice to what would have happened, if one had made a different choice in the same state of nature. There is no particular reason, insofar as one is ultimately concerned with how things actually turn out, to privilege this counterfactual outcome over all the other possible unrealized outcomes of the gambles under consideration (which are all equally non-actual).
@@ 3 Personal Good
How should the relation of personal betterness on individual lives connect to the relation of impersonal betterness on outcomes and lotteries? The following principle articulates a weak constraint:
@@
If the same individuals exist in outcomes $o_{1}$ and $o_{2}$, and $o_{1}$ is personally at least as good as $o_{2}$ for every individual and personally better for at least one individual, then $o_{1}$ is impersonally better than $o_{2}$. [^14] This principle is most plausible on the assumption that there are no sources of impersonal value in the world besides the personal value of individual lives. Pareto could reasonably be rejected, for instance, by someone who believes that natural beauty is a source of impersonal value that can outweigh differences in welfare or other kinds of personal value. Throughout this paper we make the broadly welfarist assumption that the personal value of individual lives is indeed the sole source of impersonal value. Those who reject that assumption may wish to understand our arguments as concerning not impersonal value but some narrower concept like “welfarist value”. This understanding of our arguments could also sidestep egalitarian and other “holistic” worries about Personal Good, a topic to which we’ll come shortly. (Thomas (2023: §3.1) makes a similar point in defending his principle of “Person-Based Choice” against egalitarian objections.)
This principle only concerns outcomes. How might it be extended to lotteries? We will work with the following weak extension:
@@
Suppose every possible individual $i$ has the same probability of existence in lottery $L_{1}$ as in lottery $L_{2}$. Then, (i) if for every individual $i$ and life $l$, $L_{1}$ gives $i$ at least as great a probability of a life at least as personally good as $l$, compared to $L_{2}$, then $L_{1}$ is at least as impersonally good as $L_{2}$; and (ii) if in addition, for some $i$ and $l$, $L_{1}$ gives $i$ a strictly greater probability of a life at least as personally good as $l$, then $L_{1}$ is impersonally better than $L_{2}$. [^15] Our arguments below will rely mainly on part (ii) of Personal Good, appealing to part (i) only in §6. Personal Good (especially part (i)) is closely related to the principle that McCarthy et al. (2020) call “Anteriority”.
@@
This principle extends Pareto to situations of risk, and to situations where different individuals may exist in different outcomes. But it is very modest in both respects. First, it only covers situations where every individual has a fixed probability of existence, saying nothing about choices that affect that probability for even a single individual. This means that it is compatible with a wide range of views about how to value nonexistence, including the view that existence and nonexistence are always incomparable. Second, within that restricted setting, Personal Good reflects only a minimal claim about the conditions under which one lottery is better than another from a personal perspective: $L_{1}$ is at least as good as $L_{2}$ from an individual’s perspective if the personal lottery (probability distribution over lives) that she receives from $L_{1}$ weakly stochastically dominates the personal lottery she receives from $L_{2}$; and it is strictly better if the stochastic dominance relation is strict. As described above, Stochastic Dominance is a minimal and compelling principle for ranking risky alternatives, which is compatible with a very wide range of attitudes toward risk.
@@
That said, there is precedent for rejecting Personal Good. In particular, some egalitarians will reject Personal Good because, while it allows concern for ex ante equality (equality between personal lotteries), it is incompatible with natural ways of spelling out an independent concern for ex post equality (equality between individual lives in a given outcome) (Myerson 1981). For instance, consider the following pair of lotteries:
(a) Lottery 1
Heads Tails
$P_{1}$ $1$ 00
$P_{2}$ $1$ 00
(b) Lottery 2
Heads Tails
$P_{1}$ $1$ 00
$P_{2}$ 00 $1$
Personal Good implies that these two lotteries are equally good. But since Lottery 1 guarantees perfect ex post equality, while Lottery 2 guarantees some ex post inequality, one might judge that Lottery 1 is better. Relatedly, someone who places value on ex post diversity might question Personal Good: If we replace “1” and “0” in the table above with two lives that are different but equally good, one might judge that Lottery 2 is better since it guarantees diversity while Lottery 1 guarantees bland uniformity.
We will set these objections aside, however, because (as we’ll explain in the next section) our impossibility results can be established using examples where the lotteries under comparison are intuitively equivalent in terms of both ex post equality and ex post diversity. Thus the central argument of this paper, which is built on those results, could be run using a restricted version of Personal Good that only applies when neither equality nor diversity is at stake.
@@ 4 Incomparable lives
Our arguments will require some auxiliary premises, which we now introduce. We first make an assumption that we think of as basically a convenience, without significant conceptual import:
@@
For any finite set of possible individuals and any function from those individuals to lives other than non-existence, there is an outcome where exactly those individuals exist, with exactly those lives. Moreover, for any finite set of outcomes and any probability distribution defined on that set, there is a lottery which assigns those probabilities to those outcomes.
We are assuming (as we said earlier) that the set of possible individuals is countably infinite, and that the impersonal betterness relation is defined over the whole set of lotteries. We won’t need anything like the full strength of Richness; all we need is that the domain of the betterness relation includes outcomes with the structure of our examples below.
@@
Suppose that outcomes $o_{1}$ and $o_{2}$ are identical except that one possible individual receives life $a$ in $o_{1}$ and life $b$ in $o_{2}$, where $a$ and $b$ are incomparable in terms of personal value. Then $o_{1}$ and $o_{2}$ are incomparable in terms of impersonal value.
This principle is a natural counterpart of Pareto in the context of personal incomparability: We compare outcomes by comparing them from the vantage point of each individual, and when only one individual’s fate differs between two outcomes, the impersonal betterness ordering should reflect that individual’s personal betterness ordering. When the two lives that individual might receive are personally incomparable, their incomparability is “transmitted” to the outcomes in which they figure. [^16] Like Personal Good, Incomparability Transmission is open to objections related to the impersonal value of equality and diversity. We consider those objections later in this section.
Finally, we will assume the following principle about the structure of incomparability on outcomes:
@@
Suppose that outcome $o_{1}$ is impersonally incomparable with outcomes $o_{2}$ and $o_{3}$. Then there is either
@@
    an outcome $o_{1}^{+}$ which is better than $o_{1}$ for some individual, and equally good for everyone else, while still being impersonally incomparable with both $o_{2}$ and $o_{3}$, or
@@
    an outcome $o_{1}^{-}$ which is worse than $o_{1}$ for some individual, and equally good for everyone else, while still being impersonally incomparable with both $o_{2}$ and $o_{3}$.
Insensitivity either to small improvements or to small worsenings is a paradigmatic feature of incomparability, by contrast to equal goodness. Thus it should not be controversial that, given incomparable $o_{1}$ and $o_{2}$, there is either an $o_{1}^{+}$ that improves $o_{1}$ and is still incomparable with $o_{2}$, or an $o_{1}^{-}$ that worsens $o_{1}$ and is still incomparable with $o_{2}$. The above principle adds to this uncontroversial claim two further ideas. First, it adds that we can achieve this improvement (worsening) by improving (worsening) the life of some individual in $o_{1}$. This is slightly stronger, but we do not expect it to be particularly controversial (especially given our broadly welfarist assumption that the only evaluatively significant features of an outcome are which individuals exist and what lives they lead, see n. 14). Second, it adds that when $o_{1}$ is incomparable to two further outcomes $o_{2}$ and $o_{3}$, there is a single improvement (or worsening) which preserves incomparability with both $o_{2}$ and $o_{3}$. For instance, if an outcome where Patricia is an artist is incomparable with an outcome where she is a banker and an outcome where she is a CIA operative, then we can either slightly improve or slightly worsen the first outcome (e.g., by slightly adjusting Patricia’s income) while preserving both incomparabilities. Like the first, this second claim seems plausible to us, and holds in simple, natural models of value incomparability. In any case, as described below, our argument makes only limited use of Small Improvements, and does not require it to hold in full generality.
@@
Proposition 1.
Given Richness, the principles of Negative Dominance, Personal Good, Incomparability Transmission, and Small Improvements rule out personal incomparability between lives.
This can be proved using the example from the introduction, restated for ease of reference in Table 3.
@@
Again, we have two lotteries affecting two possible individuals, $P_{1}$ and $P_{2}$. (We can assume that everyone else is assured of non-existence.) Given two incomparable lives $a$ and $b$, Incomparability Transmission implies that the outcomes $(a,a)$ and $(b,b)$ are both incomparable with the outcomes $(a,b)$ and $(b,a)$. Small Improvements implies that we can find an outcome that is equally good for one individual and better for the other than $(a,b)$ while still being incomparable with both $(a,a)$ and $(b,b)$, or else an outcome that is equally good for one individual and worse for the other, while still being incomparable with both $(a,a)$ and $(b,b)$. Let’s concentrate on the first case, and assume the improvement is to the first individual, giving this person the life $a^{+}$. (The other cases are very similar.) The outcome $(a^{+},b)$ is by assumption still incomparable with $(a,a)$ and $(b,b)$. Richness then guarantees that the domain of the impersonal betterness relation includes the two lotteries described in Table 3. Both possible outcomes of Lottery 2 are incomparable with both possible outcomes of Lottery 1, so Negative Dominance implies that Lottery 2 is not better than Lottery 1. But Lottery 2 gives $P_{1}$ a stochastically dominant personal lottery, and $P_{2}$ a stochastically equivalent lottery, so Personal Good implies that Lottery 2 is better than Lottery 1.171717Nebel (2020) also presents an opaque sweetening-inspired puzzle in a population-ethical context. His puzzle compares the outcomes $(a,b)$ and $(b^{+},a^{+})$, where $a$ and $a^{+}$ are both incomparable with $b$ and $b^{+}$. He points out that the latter outcome is intuitively better than the former, and that this verdict can be derived from Pareto, anonymity, and transitivity, but that this conflicts with the “Person-Affecting Restriction” (PAR), which claims that one outcome is better than another only if there is someone for whom it’s better. Nebel sees this as a reason to give up PAR, and we agree. But giving up PAR does not resolve our puzzle, which uses only the much weaker and more plausible principle of Incomparability Transmission. (Unlike PAR, Incomparability Transmission is consistent with the other principles Nebel invokes even in the presence of personal incomparability.) As Nebel does to Hare’s argument, it’s also possible to generate a structural analogue of our argument that inverts the roles of states and individuals. In this inverted argument, the analogue of Negative Dominance asserts that if no individual’s personal lottery under $L_{1}$ is better than any individual’s personal lottery under $L_{2}$, then $L_{1}$ is not better than $L_{2}$, while the analogue of Personal Good asserts that if in at least one state, the outcome of $L_{1}$ is Suppes-Sen superior to the outcome of $L_{2}$ (i.e., Pareto superior to a permutation of that outcome), and in all other states, the outcomes of $L_{1}$ and $L_{2}$ are identical, then $L_{1}$ is better than $L_{2}$. (To put these principles into conflict, just switch the rows and columns of Lottery 1 in Table 3, with $b$’s in the top row and $a$’s in the bottom. The argument will of course require auxiliary assumptions analogous to those used in Proposition 1. Note also that the second principle requires the introduction of states of nature, which the principles in the main text avoid. The identification of states of nature across different lotteries is analogous to the identification of individuals across different lotteries required by our principle of Personal Good.)
This argument establishes, we think, that Negative Dominance and Personal Good are jointly incompatible with the existence of incomparable lives. To briefly consider the other ways out: Richness is only used to secure the existence of two lotteries with the structure of Lottery 1 and Lottery 2 in Table 3, which are hardly exotic. Our use of Small Improvements should also be uncontroversial: given that $(a,b)$ is incomparable with $(a,a)$ and $(b,b)$, then it seems (in the spirit of insensitivity to small improvements) that it should be possible to slightly improve or worsen someone’s life in $(a,b)$ while maintaining that incomparability. [^18] Recall from n. 4 that we can’t judge $(a^{+},b)$ to be better than $(a,a)$ on the grounds that it is better for someone and worse for no one, on pain of generating cycles of betterness. We think this principle is plausible in full generality, but, as with Richness, rejecting it in particular cases would not be enough to block our argument, which only requires there to be some case with the structure of the one above.
The most debatable of our auxiliary assumptions is Incomparability Transmission. This was used to secure the conclusion that, if $a$ and $b$ are personally incomparable, then $(a,a)$ and $(b,b)$ are impersonally incomparable with $(a,b)$ and $(b,a)$. This claim seems quite plausible on its face. But as with Personal Good, it might be questioned by someone who places intrinsic value on ex post equality and/or diversity. [^19] Thanks to two anonymous reviewers for this point. As we saw in §3, someone who values equality or diversity might reject Personal Good because it takes no account of correlations between individual prospects and hence is insensitive to holistic features of outcomes. Likewise, they might reject Incomparability Transmission on the grounds that, even if lives $a$ and $b$ are incomparable from an individual perspective, one of them might yield an overall outcome that is more equal or more diverse, and therefore impersonally better. For instance, someone who values ex post diversity might claim that $(a,b)$ and $(b,a)$ are both impersonally better than $(a,a)$ and $(b,b)$, contra Incomparability Transmission.
But such equality- and diversity-based objections to our argument (whether directed at Personal Good, Incomparability Transmission, or both) can be defused by slightly modifying our central example, as illustrated in Table 4:
(a) Lottery 1
Heads Tails
$P_{1}$ $a$ $c$
$P_{2}$ $b$ $d$
(b) Lottery 2
Heads Tails
$P_{1}$ $a^{+}$ $c$
$P_{2}$ $d$ $b$
Table 4:
In this new example, there are four pairwise incomparable lives, with $P_{1}$ facing a lottery between $a$ and $c$, and $P_{2}$ facing a lottery between $b$ and $d$. As before, Lottery 2 reverses the correlations between the two personal lotteries while sweetening life $a$ for $P_{1}$. But now there is no longer any apparent difference between the possible outcomes of the two lotteries in terms of either equality or diversity. [^20] One might question whether any two outcomes consisting of two incomparable lives must be equivalent in terms of equality and diversity. But again, we don’t need this to be true in full generality. And it seems very plausible that, assuming personal incompleteness and given systematic theories of equality and diversity under incompleteness, we can find some quadruple of incomparable lives that would make the outcomes in Table 4 equivalently equal and diverse. Using this example, then, our argument would go through even if we restricted both Personal Good and Incomparability Transmission to contexts where neither equality nor diversity is at stake. We haven’t made this example the centerpiece of our discussion because it requires assuming the existence of four pairwise incomparable lives, which is not guaranteed by personal incompleteness per se. But most natural incomplete theories of personal betterness will satisfy this assumption. For instance, it would be strange to think that artist and banker are incomparable but to deny that we can find two more lives (say, CIA operative and dairy farmer) to complete the required quartet.
A notable special case of the preceding result is where one of the incomparable “lives” is nonexistence. [^21] Nonexistence incomparability is one form of incomplete personal betterness that might not allow us to find four pairwise incomparable lives: one might think that existence is incomparable with nonexistence, but that any two lives other than nonexistence can be compared. But in this case, worries about equality and diversity don’t arise for our initial argument: If one of $a$ or $b$ is nonexistence, then none of the four outcomes in the original example involve any ex post inequality or diversity. Many philosophers accept the following principle:
@@
There is some life $a$ such that, whenever two outcomes $o_{1}$ and $o_{2}$ are identical except that one possible individual does not exist in $o_{1}$ and exists with life $a$ in $o_{2}$, $o_{1}$ and $o_{2}$ are incomparable.
One motivation for this principle is the belief that existence and non-existence are incomparable in terms of personal value, coupled with a principle like Incomparability Transmission. This might motivate the view that populations of different size are always incomparable (Bader 2022a; 2022b). Alternatively, whatever one thinks about the personal value of nonexistence, one might accept Variable-Population Incompleteness to avoid undesirable conclusions about impersonal value. For instance, critical-range utilitarianism, which holds that adding an individual with slightly positive welfare to a population results in incomparability, can avoid both the Repugnant Conclusion (Parfit 1984: §131, 387–90) and the Sadistic Conclusion (Arrhenius 2000). [^22] Gustafsson (2020) proposes a version of critical-range utilitarianism that is motivated by a fourth category of value called “undistinguishedness” and shows that it escapes versions of the Repugnant and Sadistic Conclusions. But the preceding argument shows us that both these views are in tension with Negative Dominance and Personal Good, taken together:
Proposition 2.
@@
To see this, let $b$ in Table 3 represent nonexistence and $a$ be a life of the sort described by Variable-Population Incompleteness, such that adding someone to the population with life $a$ yields incomparability. Then Variable-Population Incompleteness implies that $(a,a)$ and $(b,b)$ are incomparable with $(a,b)$ and $(b,a)$. [^23] While the argument as stated appeals to the “empty population” $(b,b)$, we can avoid this by simply adding another person living life $c$ to all outcomes under consideration. The rest of the proof then follows the proof of Proposition 1. (Note that this result does not use Incomparability Transmission.) The upshot is that Negative Dominance and Personal Good together tell strongly against views that see changes in population size as a source of incomparability.